OpenStudy (anonymous):

how to write an equation for three consecutive odd numbers that equal 117???

OpenStudy (calculusxy):

What this means is x+(x+2)+(x+2)=117.

OpenStudy (anonymous):

@calculusxy Close. It's two numbers. It's \(x + (x+2) + (x+4) = 117\)

OpenStudy (calculusxy):

Since it states odd numbers, whereas x=1 and then you add two because odd numbers goes on like this.

OpenStudy (calculusxy):

But since it is consecutive, it means "one after the other".

OpenStudy (calculusxy):

@SACAPUNTAS

OpenStudy (anonymous):

Yes. Two consecutive. So for example: 1, 3, 5. 3 = 1 + 2. 5 = 1 + 4.

OpenStudy (calculusxy):